OpenStudy (anonymous):

Integrate:

OpenStudy (anonymous):

\[\large \int\limits_{}^{}e^{x^{4}}(x+x^3+2x^5)e^{x^{2}}dx\]

OpenStudy (unklerhaukus):

is this an indefinite integral ?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

then the error function isnt going to work

OpenStudy (unklerhaukus):

can you simplfy the integrand?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

try by parts separately or use this | e^x (f(x)+f`(x)) = e^x(f(x))

OpenStudy (anonymous):

what's f(x) here?

OpenStudy (anonymous):

this is what you have to find.... :D

OpenStudy (experimentx):

that you got to find int e^(g(x)) (g'(x)f(x)+f'(x)) = e^(g(x))f(x)

OpenStudy (experimentx):

put g(x) = x^4 + x^2 and g'(x) f(x) + f'(x) = 2x^5+x^3+x, put the value of g'(x) and find f(x)

OpenStudy (anonymous):

i'm not getting pls you solve

OpenStudy (experimentx):

look for the particular solution of f(x) that is x^2/2 that should give you your answer is e^(g(x)) x^2/2

OpenStudy (anonymous):

we have not studied diff. eq. ,so isn't there any other way to solve?

OpenStudy (experimentx):

this is just integration by parts. if you manipulate things carefully you should get the solution. best is to guess your solution that your solution will be of the form e^(x^4+x^2) (ax^2+bx+c)

OpenStudy (experimentx):

differentiate it and compare values and solve for a,b,c ... you can guess that your solution will be of the form because the polynomial term is or order ^5

OpenStudy (experimentx):

if it were less than x^3, i.e. e^(x^4+x^2) (x^2 + x + c) then an educated guess would be that it wouldn't be integrable in elementary terms.